# Beyond Memorization: Evaluating the True Type Inference Capabilities of LLMs for Java Code Snippets

### v1

@@ 8.2. Type Inference for Dynamically Typed Languages
Type inference for dynamically typed languages faces extra challenges as there can be variables with insufficient static type constraints in programs written in these languages, which static type inference cannot soundly handle. Many existing studies resort to deep learning-based approaches to overcome this challenge. For example, Hellendoorn et al. (2018) propose DeepTyper, a deep learning model that is trained to provide type suggestions given certain contexts and relations. Pradel et al. (2020) propose TypeWriter, which uses a deep learning-based approach to predict the types and then utilizes a search-based approach to validate the predicted type. Peng et al. (2022) propose HiTyper, which is a deep learning-based approach combining static type inference. It conducts static inference and DL-based prediction iteratively to construct a type dependency graph, which records type dependency information among variables. This paper focuses on the performance of LLMs in type inference for Java code snippets, leaving the investigation of type inference for dynamically typed languages to future work.
@@ 8.3. LLM Data Leakage
Data leakage is not limited to type inference but also affects other software engineering tasks, such as automated program repair using LLMs (Xia et al., 2023; Ouyang et al., 2024). Xia et al. (2023) discovered that 15% of the bug-fixing patches for automated program repair generated by earlier LLMs (i.e., CodeT5, GPT-Neo, GPT-J, and GPT-NeoX) were already present in the training data. Sainz et al. examined several academic datasets and reported that most of them were either used as training data for ChatGPT or likely exposed to it (Sainz et al., 2023). A subsequent work done by Golchin and Surdeanu (2023) proposes a more advanced detection technique and detects the presence of test data of several datasets in the training data of LLMs. As LLM training datasets grow larger and increasingly opaque, detecting instances of data leakage during evaluation becomes progressively more challenging.
The data leakage problem remains an open challenge for evaluating LLMs. Some studies have been conducted in other fields (Magar and Schwartz, 2022; Balloccu et al., 2024; Mirzadeh et al., 2024; Cao et al., 2024) regarding data leakage and LLM evaluation. Mirzadeh et al. (2024) showed that LLMs have significant limitations when conducting genuine mathematical reasoning which suggests that LLMs are conducting sophisticated pattern matching rather than true logical reasoning. A recent study (Cao et al., 2024) demonstrated that the performance of LLMs is far behind undergraduate students on the proposed project-level Java benchmark that exercises OOP features. Our evaluation on type inference rather than math questions or code generation also showed a significant drop in performance with generated code snippets which suggests that understanding the type system in code is challenging for LLMs. So far, there is no conclusive evidence showing that LLMs are conducting true type inference on code snippets. Small changes to the inputs can drastically alter model outputs (Shi et al., 2023; Mirzadeh et al., 2024; Jiang et al., 2024). While some guidelines have been raised for software engineering research using LLMs (Ozkaya, 2023; Sallou et al., 2024), protecting against data leakage remains an open challenge. We hope our work will shed further light on the data leakage issue in software engineering.
@@ 9. Conclusion
This paper conducted a comprehensive assessment of LLMs’ type inference capabilities on Java code snippets. Our three-pronged evaluation mitigated potential data leakage issues and identified possible limitations of LLMs. First, we create a new dataset named ThaliaType. By evaluating the performance of LLMs on ThaliaType and StatType-SO, we conclude that LLM type inference performance may not generalize well to unseen code snippets with up to 59% decrease in precision and up to 72% decrease in recall. In addition, LLM type inference performance diminishes when inferring less commonly used FQNs, potentially introducing bias and presenting opportunities for future research to enhance performance. Second, we designed and applied three semantic-preserving code transformations to code snippets in ThaliaType and StatType-SO to investigate LLMs’ semantic understanding. Through evaluating LLMs with these transformed code snippets, we find that the precision and recall of LLMs in type inference depend on the exact content of the code snippet, even when the semantic meaning of the code remains unchanged. This suggests that LLMs may rely heavily on superficial patterns in the code rather than truly understanding its semantics, harming LLMs’ performances’ generaliability. Third, we apply delta debugging to analyze the minimal syntactic elements sufficient for LLMs to infer types. We find that LLMs can still correctly infer FQNs even when the types were absent from the snippets, that StatType-SO may allow LLMs to rely on superficial patterns rather than deep code understanding. Our experiments suggest that LLMs’ strong performance in prior evaluations was likely due to data leakage rather than a genuine understanding and analysis of the semantics in code snippets. Our findings suggest that future evaluations of LLMs should utilize a generated dataset, carefully designed to assess the true capabilities of LLMs for type inference tasks. All of our code and dataset are available at https://github.com/uw-pluverse/thalia-type.

### v2

@@ 8.2. Type Inference for Dynamically Typed Languages
Type inference for dynamically typed languages faces extra challenges as there can be variables with insufficient static type constraints in programs written in these languages, which static type inference cannot soundly handle. Many existing studies resort to deep learning-based approaches to overcome this challenge. For example, Hellendoorn et al. (2018) propose DeepTyper, a deep learning model that is trained to provide type suggestions given certain contexts and relations. Pradel et al. (2020) propose TypeWriter, which uses a deep learning-based approach to predict the types and then utilizes a search-based approach to validate the predicted types. Peng et al. (2022) propose HiTyper, a deep learning-based approach that combines with static type inference. It conducts static inference and DL-based prediction iteratively to construct a type dependency graph, which records type dependency information among variables. This paper focuses on the performance of LLMs in type inference for Java code snippets, leaving the investigation of type inference for dynamically typed languages to future work.
@@ 8.3. LLM Data Leakage
Data leakage is not limited to type inference but also affects other software engineering tasks. Currently, LLMs have been integrated into a wide variety of software engineering tools such as automated code repair (Yuan et al., 2022; Zhang et al., 2024; Xia et al., 2023; Xia and Zhang, 2024; Ouyang et al., 2024; Bouzenia et al., 2025; Ye et al., 2025; Kong et al., 2025b), code generation (Jiang et al., 2024a; Mu et al., 2024; Li et al., 2025; Fan et al., 2025), code refactoring (Shirafuji et al., 2023; Pomian et al., 2024), and code completion (Guo et al., 2023; Wu et al., 2024; Semenkin et al., 2025), which can all be potentially affected by data leakage either currently or in the future.
Xia et al. (2023) discovered that 15% of the bug-fixing patches for automated program repair generated by earlier LLMs (i.e., CodeT5, GPT-Neo, GPT-J, and GPT-NeoX) were already present in the training data. Sainz et al. (2023b, a) examined several academic datasets and reported that most of them were either used as training data for ChatGPT or likely exposed to it. A subsequent work by Golchin and Surdeanu (2023) proposes a more advanced detection technique and detects the presence of test data of several datasets in the training data of LLMs. Marone and Van Durme (2023) propose data portraits to efficiently check whether a given string is present in the training data, assuming the training data is available. Recently, Kong et al. (2025a), detected memorization in LLMs using low-probability events such as exact code repair matches. However, for type inference on code snippets, there is generally one ground truth of types that are expected. Thus, exact matches do not precisely indicate data leakage. As LLM training datasets grow larger and increasingly opaque, detecting instances of data leakage during evaluation becomes progressively more challenging.
Nevertheless, data leakage remains an open challenge for evaluating LLMs. Some studies have been conducted in other fields (Magar and Schwartz, 2022; Balloccu et al., 2024; Mirzadeh et al., 2024; Cao et al., 2024; Uddin et al., 2025; Elangovan et al., 2021) regarding data leakage and LLM evaluation. Mirzadeh et al. (2024) showed that LLMs have significant limitations when conducting genuine mathematical reasoning, which suggests that LLMs are conducting sophisticated pattern matching rather than true logical reasoning. A recent study by (Cao et al., 2024) demonstrated that the performance of LLMs is far behind undergraduate students on the proposed project-level Java benchmark that exercises object-oriented programming features.
Our evaluation on type inference rather than math questions or code generation also showed a significant drop in performance with generated code snippets which suggests that understanding the type system in code is challenging for LLMs. So far, there is no conclusive evidence showing that LLMs are conducting genuine type inference on code snippets. Small changes to the inputs can drastically alter model outputs (Shi et al., 2023; Mirzadeh et al., 2024; Jiang et al., 2024b). While some guidelines have been raised for software engineering research using LLMs (Ozkaya, 2023; Sallou et al., 2024), protecting against data leakage remains an open challenge. We hope our work will shed further light on the data leakage issue in software engineering.
@@ 9. Conclusion
This paper conducted a comprehensive assessment of LLMs’ type inference capabilities on Java code snippets. First, to address potential data leakage we introduced a new benchmark suite, ThaliaType, and compared it against the widely used StatType-SO. Second, using StarCoder2 as a baseline (which we confirmed to suffer from data leakage), our evaluation revealed that all tested LLMs exhibited similar performance degradations on unseen code snippets. Specifically, we observed up to a 59% decrease in precision and a 72% decrease in recall. Moreover, LLM performance diminished when inferring less commonly used FQNs, suggesting potential bias and presenting opportunities for future research. Finally, we designed and applied three semantic-preserving code transformations to both benchmark suites, to investigate LLMs’ understanding of the execution semantics. While LLMs maintained consistent performance under simple transformations, their performance significantly declined when transformations are combined on StatType-SO. These results on StatType-SO, not observed on ThaliaType, indicate that LLM performance on StatType-SO may be inflated by data leakage and may not be generalizable. Overall, our findings underscore the need for future evaluations to incorporate unseen benchmarks, such as newly generated ThaliaType code snippets and transformed variants, while discontinuing StatType-SO. This approach provides a more reliable assessment of LLMs’ performance in type inference tasks. All code and datasets used in this paper are publicly available at https://github.com/uw-pluverse/thalia-type.